Suppose $500 is invested at 6% annual inerest compounded twice a year. When will the investment be worth $1000?
Answer by stanbon(48558)

Suppose $500 is invested at 6% annual inerest compounded twice a year. When will the investment be worth $1000?
----------------
A(t) = P(1+(r/n))^(nt)
--------------------------
1000 = 500(1+(0.06/2))^(2t)
2 = (1.03)^2t
2t*log(1.03) = log2
2t = log2/log1.03
2t = 23.4498
t = 11.7249 years
======================
Cheers,
Stan H.